# Integrated modeling of boron powder injection for real-time plasma-facing component conditioning

### v2

@@ Abstract
An integrated modeling framework for investigating the application of solid boron (B) powder injection for real-time surface conditioning of plasma-facing components (PFCs) in tokamak environments is presented. Utilizing the DIII-D impurity powder dropper (IPD) setup, this study simulates B powder injection scenarios ranging from milligrams to tens of milligrams per second, corresponding to boron flux rates of $10^{20}-10^{21}$ B/s in standard L-mode conditions. The comprehensive modeling approach combines EMC3-EIRENE for simulating the deuterium plasma background and the Dust Injection Simulator (DIS) for the ablation and transport of the boron powder particles. The results show substantial transport of boron to the inboard lower divertor, predominantly influenced by the main ion plasma flow. The dependency on powder particle size (5-250 $\mu$m) was found to be insignificant for the scenario considered. The effects of erosion and redeposition were considered to reconcile the discrepancies with experimental observations, which saw substantial deposition on the outer divertor plasma-facing components. For this purpose, the WallDYN3D code was updated to include boron sources within the plasma domain and integrated into the modeling framework. The mixed-material migration modeling shows evolving boron deposition patterns, suggesting the formation of mixed B-C layers or predominantly B coverage depending on the powder mass flow rate. While the modeling outcomes at lower B injection rates tend to align with DIII-D experimental observations, the prediction of near-pure boron layers at higher rates has yet to be experimentally verified in the carbon environment of the DIII-D tokamak. The extensive reach of boron layers found in the modeling suggests the need for modeling that encompasses the entire wall geometry for more accurate experimental correlations. This integrated approach sets a precedent for analyzing and applying real-time in-situ boron coating techniques in advanced tokamak scenarios, potentially extendable to ITER.
@@ 1 Introduction
This paper presents a new framework for modeling solid boron injection, the key material for real-time wall conditioning in tokamaks, stellarators, and heliotrons. It integrates transport, erosion, and deposition modeling to assess boron coverage on divertor PFCs, crucial for evaluating its suitability in future devices like ITER and FPPs [14, 15, 16]. The goal is to demonstrate proof of concept by coupling three models based on a DIII-D RTWC scenario from a prior study [17].
@@ 2 Coupled dust and plasma impurity transport modeling with DIS and EMC3-EIRENE
where $R_{f}$ describes the force acting on the powder particles consisting of the drag force exerted by the plasma ($F_{drag}$), electrostatic force ($F_{E}$), gravitational force ($F_{grav}$), and centrifugal force ($F_{centr.}$). Presently, $F_{E}$ does not contribute to the dust transport since EMC3-EIRENE does not yet solve the electric field self-consistently. Particles reaching the wall collide elastically, reflect back, and continue evaporating in the plasma. The deposition depths in the plasma of the injected materials are determined by material-specific properties such as vapor pressure, heat sublimation, emissivity, secondary electron emission, material density, and boiling and melting points. Note that the present version of the model assumes perfect spherical symmetry of particles, neglecting effects such as the rocket force [28].
@@
where $S_{z}$ is the impurity ionization source for charge state $Z$, $D_{imp}$ the anomalous diffusivity, $n_{z}$ and $p_{z}$ the impurity density and pressure, $E_{\parallel}$ the parallel electric field, and $V_{\parallel}$ and $V_{z\parallel}$ the main and impurity ion parallel velocities. The impurity-main ion collision time is given by $\tau_{zi}=\frac{1.47\cdot 10^{13}m_{z}(\nicefrac{{T_{i}^{3}}}{{m_{i}}})^{0.5}}{(1+\nicefrac{{m_{i}}}{{m_{z}}})n_{i}Z^{2}ln\Lambda}$ with $\ln\Lambda\approx 15$ [29]. Equation 3 terms correspond to pressure ($F_{P}$), electric ($F_{E}$), friction ($F_{fr}$), and electron/ion thermal forces ($F_{th,e}$, $F_{th,i}$). The impurity temperature is assumed equal to the main ion temperature ($T_{z}=T_{i}$).
In many plasma edge scenarios, electrostatic and pressure forces can be neglected, reducing equation 3 to focus on friction and ion thermal forces, independent of charge state $Z$. At steady state, this leads to a flow velocity balance [30, 17]:
@@ 3 Coupling impurity transport with erosion and deposition dynamics in WallDYN3D
$\frac{d\delta_{ei,wk}}{dt}=\sum_{qi}\Gamma_{ei,wj,qi}^{In}-(\Gamma_{wk,qi}^{Ero}+\Gamma_{ei,wk}^{Refl})+\Gamma_{wk,qi}^{Bulk},$ (6)
where $\Gamma_{wk,qi}^{Bulk}$ represents a bulk material reservoir, that maintains the areal density of all wall elements.
Figure 1: (a) Top-down view and (b) cross-sectional view on boron source distribution resulting from ablation of 5-250 $\mu$m size powder particles along the trajectories calculated with DIS. (c) Boron flux profile at the lower inner divertor (peak location) for 5-250 $\mu$m B powder particles (zero impurity recycling), shown along the Z-coordinate of the inner divertor target.
@@ 4 Full-torus modeling of boron powder injection and transport with EMC3-EIRENE and DIS
Figure 1(a,b) displays boron source distributions along particle trajectories for sizes from 5 to 250 $\mu$m, representing typical experimental diameters. The injected mass flow rate is constant for all particle sizes. The evaporation usually occurs when the powder particles reach the separatrix, where the electron temperature and associated heat flux $Q_{subl}$ increase, raising the powder particle temperature to the evaporation point. For the mass loss rate follows: $\frac{dm_{d}}{dt}\propto Q_{subl}$. Larger particles, such as those measuring 150 and 250 $\mu$m, follow a predominantly vertical trajectory due to gravity and may cross the separatrix before fully ablating. Conversely, the trajectories of smaller particles (5 $\mu$m) are primarily influenced by the drag and centrifugal forces exerted by the toroidal plasma flow, which pushes these unconfined particles outward. All particle sizes considered in this study evaporate within the simulation domain, with boron ions crossing the inner boundary handled by the 1D core impurity model [18].
@@
A notable increase in boron density is found on the high-field side of the plasma. This inboard peaking of the B density is attributed to the dominant parallel friction force terms associated with the parallel main ion flow, as delineated in equations 2 and 3 and illustrated as Mach flow profile in Figure 2. The parallel main ion flow in the powder deposition region shows a dominant component toward the inboard target (blue region), i.e., pushing the boron impurities effectively poloidally counter-clockwise. Thus, the inward directional bias of the flow culminates in heightened boron densities and fluxes impacting the plasma-facing components on the high-field side.
@@
The boron incident flux distributions on the inner and outer divertor targets reveal a substantial asymmetry, with over $95\%$ of the total boron flux directed towards the inboard target according to the trace impurity fluid model. The toroidally averaged boron fluxes are shown for all particle sizes in Figure 1(c). The particle size does not significantly affect this inboard/outboard flux imbalance, with the frictional forces ($\propto V_{fr}\propto M_{||}$) of the main ion flow primarily governing the transport of boron in the plasma scenario considered. However, this boron distribution only represents direct deposition from the injection source without considering any subsequent removal or migration upon contact with the target surfaces, i.e., impurities do not recycle. As a result, deposition is narrowly confined to the far SOL, specifically between $Z=-115$ and $-95\,\text{cm}$ on the inboard divertor. Additionally, toroidal asymmetries in the boron deposition on the outer divertor arise from the localized nature of the boron source and the specific orientation of the flux tubes connecting the source region to the divertor targets, as noted in [17]. A limitation of this scenario is that the plasma simulation domain covers only the divertor region, as the plasma grid was originally designed for transport studies rather than plasma-material interaction.
@@
Figure 4: (a) Visualization of the lower divertor region with wall indices (IDX) representing poloidal wall elements. Blue elements are within the simulation domain. Charge state-integrated 2D boron redistribution matrices, as toroidal averages of the 3D matrix, are shown for (b) the fluid model and (c) the kinetic model. ”Source wall IDX” indicates where impurities are eroded, while ”Destination wall IDX” shows where they are deposited.
@@
A 3D visualization of the boron fluxes impacting the divertor targets at an IPD injection rate of $10^{20}$ B/s, as simulated using WallDYN3D, is shown in Figure 3. This figure initially depicts the boron deposition without reflections, erosion, or redeposition (Figure 3(a)). It then progresses through time steps from 0 to 400 seconds, incorporating these effects (Figure 3(b-f)). Notably, after reaching quasi-equilibrium after approximately 20 seconds, there is a significant alteration in the boron flux distribution. The outboard divertor target experiences an increased flux compared to the inboard target, as detailed in figures 3(e,f).
@@
In WallDYN3D, impurity transport is treated kinetically by default [31]. A comparison between kinetic and fluid modeling was conducted to investigate potential differences in the transport between the high and low-field sides of relatively light B particles. In this case, boron is sourced from the high field side target. The three-dimensional charge-state integrated boron redistribution matrices have been toroidally averaged, resulting in two-dimensional representations, since this scenario assumed the wall geometry was toroidally symmetric. Figure 4(a) shows the wall elements and corresponding indices (IDX) in the poloidal plane. The boron redistribution matrices are presented for both the kinetic (Figure 4b) and fluid (Figure 4c) approaches. The matrices indicate impurity erosion and deposition locations, specifically on the divertor wall elements. The diagonal terms indicate the dominating local redeposition of B on the inner ($IDX=6-12$) and outer divertor targets ($IDX=19-28$). The matrices also reveal an absence of transport from the inner to the outer target. However, values in the bottom right sections indicate additional transport from the outer to the inner target. This behavior is consistent in both the fluid and kinetic treatment qualitatively and quantitatively. The remaining differences between the kinetic and fluid models stem from how they handle impurity transport near the boron injection location. The fluid model assumes immediate equilibrium with the plasma, simplifying the process and leading to overestimated long-range transport. In contrast, the kinetic model accounts for the time needed for boron ions to accelerate and thermalize, resulting in more accurate, localized deposition.
@@ 6 Impact of boron source strength on mixed-material coating composition
The dynamics of the B-C surface composition are affected by erosion and reflection properties on wall elements. One-dimensional boron and carbon flux profiles at the toroidal DiMES location (dashed vertical line, [33]) are shown in figure 5 for time steps of 0, 1, 10, 20, and 400 seconds, with boron influx rates of $10^{20}$ and $10^{21}$ B/s (see poloidal wall index in figure 4a). The WallDYN3D calculations indicate that boron flux increases significantly due to reerosion, migration, and redeposition while carbon flux decreases. At the lower boron rate, boron flux to the high-field side target (inner peak) increases over time from $\approx 6\times 10^{19}$ to $1.2\times 10^{20}$ m-1s-1, and more dramatically on the outboard side (outer peak) from $\approx 3\times 10^{19}$ to $9\times 10^{20}$ m-1s-1. The carbon flux sees only a slight reduction. Conversely, at the higher boron rate ($10^{21}$ B/s), peak levels stabilize quickly, within 1 second at the inboard target, with carbon levels dropping by about 10%.
@@
Adjusting the boron injection rate modifies the surface composition of divertor targets, as shown in Figure 6. With a boron influx rate of $10^{20}$ B/s, the B surface concentrations on the inner and outer targets reach approximately 35%, with boron area densities ranging from $0.3-1.4\times 10^{16}$ cm-2. Increasing the influx rate tenfold to $10^{21}$ B/s significantly enriches the boron on the target surfaces, potentially forming nearly pure boron layers, with area densities between $1.8$ and $3.6\times 10^{16}$ cm-2. This saturation phenomenon contrasts with lower influx rates, where the surface composition sustains a mixed B-C coating, approximating a B concentration closer to experimental findings, with B and C concentrations each being 40% as shown in figure 6b. This effect can be attributed to the differential sputtering rates of B and C, which minimize the overall sputtering yields. This effect and associated boron-carbon compositions have been observed in previous DIII-D IPD experiments and traditional boronization procedures in tokamaks like TEXTOR and ASDEX-Upgrade[2, 34, 35]. For IPD experiments, however, achieving a balanced B:C concentration typically required B powder rates of 10-20 mg/s or $10^{21}$ B/s. Discussions in the literature, such as in [2], highlight a marginal dependence of surface composition on boron influx rates, underscoring the challenges of attaining pure boron layers in environments with significant carbon presence. Differences between modeling and experiment may be partially explained by the simulation domain focusing on the divertor region, with the main chamber wall not yet included. Carbon wall elements beyond the divertor might contribute to the total C influxes due to additional plasma-wall interactions (e.g., main wall CX-erosion) not considered in the present study.

### v3

@@ Abstract
An integrated modeling framework for investigating the application of solid boron (B) powder injection for real-time surface conditioning of plasma-facing components (PFCs) in tokamak environments is presented. Utilizing the DIII-D impurity powder dropper (IPD) setup, this study simulates B powder injection scenarios ranging from milligrams to tens of milligrams per second, corresponding to boron flux rates of $10^{20}-10^{21}$ B/s in standard L-mode conditions. The comprehensive modeling approach combines EMC3-EIRENE for simulating the deuterium plasma background and the Dust Injection Simulator (DIS) for the ablation and transport of the boron powder particles. EMC3 trace impurity fluid modeling results show substantial boron transport to the inboard lower divertor, predominantly influenced by the main ion plasma flow. The dependency on powder particle size (5-250 $\mu$m) was found to be insignificant for the scenario considered. The effects of erosion and redeposition were considered to reconcile the discrepancies with experimental observations, which saw substantial deposition on the outer divertor plasma-facing components. For this purpose, the WallDYN3D code was updated to include boron sources within the plasma domain and integrated into the modeling framework. The mixed-material migration modeling shows evolving boron deposition patterns, suggesting the formation of mixed B-C layers or predominantly B coverage depending on the powder mass flow rate. While the modeling outcomes at lower B injection rates tend to align with DIII-D experimental observations, the prediction of near-pure boron layers at higher rates has yet to be experimentally verified in the carbon environment of the DIII-D tokamak. The extensive reach of boron layers found in the modeling suggests the need for modeling that encompasses the entire wall geometry for more accurate experimental correlations. This integrated approach sets a precedent for analyzing and applying real-time in-situ boron coating techniques in advanced tokamak scenarios, potentially extendable to ITER.
@@ 1 Introduction
This paper presents a new framework for modeling solid boron injection, the key material for real-time wall conditioning in tokamaks, stellarators, and heliotrons. It integrates transport, erosion, and deposition modeling to assess boron coverage on divertor PFCs, crucial for evaluating its suitability in future devices like ITER and Fusion Pilot Plants (FPPs) [14, 15, 16]. As a proof of concept, results from the new workflow, which couples three key models — EMC3-EIRENE, DIS, and WallDYN3D — based on a DIII-D RTWC scenario from a previous study [17], will be evaluated.
@@ 2 Coupled dust and plasma impurity transport modeling with DIS and EMC3-EIRENE
where $R_{f}$ describes the force acting on the powder particles consisting of the drag force exerted by the plasma ($F_{drag}$), electrostatic force ($F_{E}$), gravitational force ($F_{grav}$), and centrifugal force ($F_{centr.}$). Presently, $F_{E}$ does not contribute to the dust transport since EMC3-EIRENE does not yet solve the electric field self-consistently. Particles reaching the wall collide elastically, reflect, and continue evaporating in the plasma. The deposition depths in the plasma of the injected materials are determined by material-specific properties such as vapor pressure, heat sublimation, emissivity, secondary electron emission, material density, and boiling and melting points. Note that the present version of the model assumes perfect spherical symmetry of particles, neglecting effects such as the rocket force [28].
@@
where $S_{z}$ is the impurity ionization source for charge state $Z$, $D_{imp}$ the anomalous diffusivity, $n_{z}$ and $p_{z}$ the impurity density and pressure, $E_{\parallel}$ the parallel electric field, and $V_{\parallel}$ and $V_{z\parallel}$ the main and impurity ion parallel velocities. The impurity-main ion collision time is given by $\tau_{zi}=\frac{1.47\cdot 10^{13}m_{z}(\nicefrac{{T_{i}^{3}}}{{m_{i}}})^{0.5}}{(1+\nicefrac{{m_{i}}}{{m_{z}}})n_{i}Z^{2}ln\Lambda}$ with $\ln\Lambda\approx 15$ [29]. Equation 3 terms correspond to pressure ($F_{P}$), electric ($F_{E}$), friction ($F_{fr}$), and electron/ion thermal forces ($F_{th,e}$, $F_{th,i}$). The impurity temperature is assumed to be equal to the main ion temperature ($T_{z}=T_{i}$).
In many plasma edge scenarios, electrostatic and pressure forces can be neglected, reducing equation 3 to focus on friction and ion thermal forces independent of charge state $Z$. At steady state, this leads to a flow velocity balance [30, 17]:
@@ 3 Coupling impurity transport with erosion and deposition dynamics in WallDYN3D
$\frac{d\delta_{ei,wk}}{dt}=\sum_{qi}\Gamma_{ei,wj,qi}^{In}-(\Gamma_{ei,wk}^{Ero}+\Gamma_{ei,wk}^{Refl})+\Gamma_{ei,wk}^{Bulk},$ (6)
where $\Gamma_{ei,wk}^{Bulk}$ represents a bulk material reservoir, that maintains the areal density of all wall elements.
Figure 1: (a) Top-down view and (b) cross-sectional view on boron source distribution resulting from ablation of 5-250 $\mu$m size powder particles along the trajectories calculated with DIS. (c) Boron flux profile at the lower inner divertor (peak location) for 5-250 $\mu$m B powder particles (zero impurity recycling), shown along the Z-coordinate of the inner divertor target. The vertical dashed line indicates the inner strike point position.
@@ 4 Full-torus modeling of boron powder injection and transport with EMC3-EIRENE and DIS
Figure 1(a,b) displays boron source distributions along particle trajectories for sizes from 5 to 250 $\mu$m, representing typical experimental diameters. The injected mass flow rate is constant for all particle sizes. The evaporation usually occurs when the powder particles reach the separatrix, where the electron temperature and associated heat flux $Q_{evap}$ increase, raising the powder particle temperature to the evaporation point. For the mass loss rate follows: $\frac{dm_{d}}{dt}\propto Q_{evap}$. Larger particles, such as those measuring 150 and 250 $\mu$m, follow a predominantly vertical trajectory due to gravity and may cross the separatrix before fully ablating. Conversely, the trajectories of smaller particles (5 $\mu$m) are primarily influenced by the drag and centrifugal forces exerted by the toroidal plasma flow, which pushes these unconfined particles outward. All particle sizes considered in this study evaporate within the simulation domain, with boron ions crossing the inner boundary handled by the 1D core impurity model [18].
@@
A notable increase in boron density is found on the high-field side of the plasma. This inboard peaking of the B density is attributed to the dominant parallel friction force terms associated with the parallel main ion flow, as delineated in equations 2 and 3 and illustrated as Mach flow profile in Figure 2(h). The parallel main ion flow in the powder deposition region shows a dominant component toward the inboard target (blue region), i.e., pushing the boron impurities effectively poloidally counter-clockwise. Thus, the inward directional bias of the flow culminates in heightened boron densities and fluxes impacting the plasma-facing components on the high-field side.
@@
The boron incident flux distributions on the inner and outer divertor targets reveal a substantial asymmetry, with over $95\%$ of the total boron flux directed towards the inboard target according to the EMC3 trace impurity fluid model. The toroidally averaged boron fluxes are shown for all particle sizes in Figure 1(c). The particle size does not significantly affect this inboard/outboard flux imbalance, with the frictional forces ($\propto V_{fr}\propto M_{||}$) of the main ion flow primarily governing the transport of boron in the plasma scenario considered. However, kinetic modeling in the following will show a reduced imbalance in the fluxes between the inner and outer targets. Moreover, this boron distribution only represents direct deposition from the injection source without considering any subsequent removal or migration upon contact with the target surfaces, i.e., impurities do not recycle. As a result, deposition is narrowly confined to the far SOL, specifically between $Z=-115$ and $-95\,\text{cm}$ on the inboard divertor. Additionally, toroidal asymmetries in the boron deposition on the outer divertor arise from the localized nature of the boron source and the specific orientation of the flux tubes connecting the source region to the divertor targets, as noted in [17]. A limitation of this scenario is that the plasma simulation domain covers only the divertor region, as the plasma grid was originally designed for transport studies rather than plasma-material interaction.
@@
[h!]
Figure 4: (a) Visualization of the lower divertor region with wall indices (IDX) representing poloidal wall elements. Blue elements are within the simulation domain. Charge state-integrated 2D boron redistribution matrices, as toroidal averages of the 3D matrix, are shown for (b) the fluid model and (c) the kinetic model. ”Source wall IDX” indicates where boron impurities are eroded, while ”Destination wall IDX” shows where they are likely deposited. Each matrix entry reflects the probability or flux fraction of boron migrating from a specific source to a deposition location.
@@
A 3D visualization of the boron fluxes impacting the divertor targets at an IPD injection rate of $10^{20}$ B/s, as simulated using WallDYN3D, is shown in Figure 3. Here, the initial boron deposition shows less imbalance between the inner and outer targets compared to the EMC3 fluid model (Figure 1(c)). In the progress through time steps from 0 to 400 seconds, reflections, erosion, and redeposition are taken into account (Figure 3(b-f)). After reaching quasi-equilibrium after approximately 20 seconds, there is a significant alteration in the boron flux distribution. The outboard divertor target experiences an increased flux compared to the inboard target, as detailed in figures 3(e,f).
@@
In WallDYN3D, impurity transport is treated kinetically by default [31]. A comparison between kinetic and fluid modeling was conducted to investigate potential differences in the long-range transport between the high and low-field sides of relatively light B particles. In this case, boron is sourced from the high field side target. The three-dimensional charge-state integrated boron redistribution matrices (see equation (5)) have been toroidally averaged, resulting in two-dimensional representations since this scenario assumed the wall geometry was toroidally symmetric. Figure 4(a) shows the wall elements and corresponding indices (IDX) in the poloidal plane. The boron redistribution matrices for both the kinetic (Figure 4b) and fluid (Figure 4c) approaches display the probability distribution of where impurities, once eroded, will travel and redeposit, specifically on the divertor wall elements. The diagonal terms indicate the dominating local redeposition of B on the inner ($IDX=6-12$) and outer divertor targets ($IDX=19-28$). The matrices also reveal an absence of transport from the inner to the outer target. However, values in the bottom right sections indicate additional transport from the outer to the inner target. On the inner target, less local redeposition occurs than on the outer target, with less eroded B (and C) returning to their origin (see diagonal elements in Figure 4b,c). This behavior is consistent in both the fluid and kinetic treatment qualitatively and quantitatively within the WallDYN3D model. The remaining differences between the kinetic and fluid models stem from how they handle impurity transport near the boron injection location. The fluid model assumes immediate equilibrium with the plasma, simplifying the process and leading to overestimated long-range transport. In contrast, the kinetic model accounts for the time needed for boron ions to accelerate and thermalize, resulting in more accurate, localized deposition.
@@ 6 Impact of boron source strength on mixed-material coating composition
The dynamics of the B-C surface composition are affected by erosion and reflection properties on wall elements. One-dimensional boron and carbon flux profiles at the toroidal DiMES location (dashed vertical line, [33]) are shown in figure 5 for time steps of 0, 1, 10, 20, and 400 seconds, with boron influx rates of $10^{20}$ and $10^{21}$ B/s (see poloidal wall index in figure 4a). The WallDYN3D calculations indicate that boron flux increases significantly due to reerosion, migration, and redeposition while carbon flux decreases. At the lower boron rate, boron flux to the high-field side target (inner peak) increases over time from $\approx 6\times 10^{19}$ to $1.2\times 10^{20}$ m-2s-1, and more dramatically on the outboard side (outer peak) from $\approx 3\times 10^{19}$ to $9\times 10^{20}$ m-2s-1. The carbon flux sees only a slight reduction. Conversely, at the higher boron rate ($10^{21}$ B/s), peak levels stabilize quickly, within 1 second at the inboard target, with carbon levels dropping by about 10%.
@@
Most of the eroded boron (B) and carbon (C) recycle in front of the target through repeated erosion and re-deposition. Local re-deposition is lower on the inner target than the outer target, as shown by the migration matrix. Consequently, less of the eroded material returns to its origin on the inner target, whereas the outer target retains more due to higher re-deposition rates. The erosion yields of B layers are higher than those of carbon, and for a 50:50 B-to-C surface, the erosion flux is greater for B. The increased B influx is primarily driven by local recycling, not long-range transport (as shown in the previous section). The influx of B on the outer target is higher simply because more erosion and reflection flux is coming back to the target. The surface composition is due to the dynamic balance of erosion vs. influx. On the outer target, high re-deposition rates result in approximately zero net erosion of C and minor accumulation of B, supported by the continuous B source from the impurity powder dropper. In contrast, the inner target exhibits net erosion of C and nearly zero net deposition of B due to lower re-deposition. As stated before, some B is lost (especially from the inner target).
Adjusting the boron injection rate modifies the surface composition of divertor targets, as shown in Figure 6. With a boron influx rate of $10^{20}$ B/s, the B surface concentrations on the inner and outer targets reach approximately 35%, with boron area densities ranging from $0.3-1.4\times 10^{16}$ cm-2. Increasing the influx rate tenfold to $10^{21}$ B/s significantly enriches the boron on the target surfaces, potentially forming nearly pure boron layers, with area densities between $1.8$ and $3.6\times 10^{16}$ cm-2. This saturation phenomenon contrasts with lower B injection rates, where the surface composition sustains a mixed B-C coating, approximating a B concentration closer to experimental findings, with B and C concentrations each being 40% as shown in figure 6(b). This effect can be attributed to the differential sputtering rates of B and C, which minimize the overall sputtering yields. It has been observed in previous DIII-D IPD experiments and traditional boronization procedures in tokamaks like TEXTOR and ASDEX-Upgrade [2, 34, 35]. For IPD experiments, however, achieving a balanced B:C concentration typically required B powder rates of 10-20 mg/s or $10^{21}$ B/s. Discussions in the literature, such as in [2], highlight a marginal dependence of surface composition on boron influx rates, underscoring the challenges of attaining pure boron layers in environments with significant carbon presence. Differences between modeling and experiment may be partially explained by the simulation domain focusing on the divertor region, with the main chamber wall not yet included. Carbon wall elements beyond the divertor might contribute to the total C influxes due to additional plasma-wall interactions (e.g., main wall CX-erosion) not considered in the present study.